# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ 2.3. Technical Lemmas on $C^{0}$-Riemannian manifolds
Combining again the fact that $M$ is $C^{0}_{\rm loc}$-asymptotic to $\mathbb{H}^{n}_{K}$ with item (1) in Section 2.3 we get that for any $R>0$ there exists a constant $C_{R}>0$ such that $C_{R}^{-1}r^{n}\leq|B_{r}(x)|\leq C_{R}r^{n}$ for any $x\in M$ and $r\in(0,R)$. Hence recalling that $(M,\mathsf{d},\mathcal{H}^{n})$ is also PI by Section 2.3, then it is well-known that there holds a relative isoperimetric inequality in balls of $M$, see [HajlaszKoskela, Theorem 5.1] and [Ambrosio02, Remark 4.4]. Since also $\inf_{x\in M}|B_{1}(x)|>0$ thanks to the asymptotic assumption, it is readily checked that the proof of [ANP, Lemma 2.10] can be repeated in our setting, yielding that: for any $R>0$ there exists $\widetilde{C}=\widetilde{C}(R)>0$ such that for any $E\subset M$ with $|E|\in(0,+\infty)$ there exists $x_{E}$ such that
@@
Fix now $\overline{V}\in(0,+\infty)$ and $\eta\in(0,\overline{V})$. Let also $V_{0}\in(\max\{\overline{V}-1,\eta\},\overline{V}+1)$ and for any $\varepsilon>0$ let $E\subset M$ be a bounded set such that $|E|=V_{0}$ and $P(E)\leq I(V_{0})+\varepsilon$ (it is readily checked that, arguing as in [AFP21, Lemma 2.17], the isoperimetric profile is achieved by bounded sets).
@@
The next theorem is based on a concentration-compactness argument that has been used several times in the literature, applied to the study of the isoperimetric problem in noncompact manifolds. In the smooth setting it has been first obtained in [Nardulli, Theorem 2]. See also [RitoreBook, Theorem 4.48] and references therein for a complete account. Results analogous to Theorem 2.2 have been worked out also in the setting of nonsmooth spaces with bounds below on the curvature, see [AFP21, Theorem 4.6] and [ANP, Theorem 3.3 & Theorem 1.1].
@@
Since the proof of Theorem 2.2 is standard and follows closely the strategy of [AFP21, ANP] we just sketch it.
At first, one gets the analogue of Ritoré–Rosales’ result [ANP, Theorem 3.3] in the setting of Theorem 2.2. Indeed, the proof of [ANP, Theorem 3.3] uses the coarea formula, the precompactness of $BV$ in $L^{1}_{\mathrm{loc}}$, and the existence, around every point $p\in M$, of a one-parameter family $\{\mathcal{B}_{p,r}\}_{r\in(0,\varepsilon)}$ of sets such that $r\mapsto|\mathcal{B}_{p,r}|$ is continuous and vanishing as $r\to 0$, and $P(\mathcal{B}_{p,r})\to 0$ as $r\to 0$. The first two come from Section 2.3, and Theorem 2.1, while for the last one it suffices to take pre-images of small Euclidean balls under the map in the second item of Section 2.1, as it has been done in the proof of Section 2.3. Once this is done, one follows verbatim the proof of [AFP21, Theorem 4.6], which additionally needs that $(M,\mathsf{d},\mathcal{H}^{n})$ is PI, and $|B_{r}(p)|\geq\zeta r^{3}$ for every $r\in(0,R]$, and every $p\in M$, where $\zeta,R>0$ are constants depending on $M$: these two properties come from Section 2.3 and from its proof (compare also with item (1) in Section 2.3).
Following [AFP21, Theorem 4.6] until (4.20), and then jumping to Step 5 in there, one finally gets the following. There is $\overline{N}\in\mathbb{N}\cup\{+\infty\}$ and radii $R_{i}\to+\infty$, $T_{i,j}\to_{i}+\infty$ for $11\leq j<\overline{N}+$ 1, and there are mutually (with respect to $j$) diverging $p_{i,j}\in M\setminus B_{R_{i}}(o)$ such that
[…]
where $B_{j}$ is a ball in $\mathbb{H}^{n}_{K}$ for any $j<\overline{N}+1$, and $E$ is an isoperimetric set in $M$. The convergence in $L^{1}$ to the $B_{j}$’s has to be intended as in the statement of Section 2.3, after the composition with biLipschitz embeddings. Moreover one has
@@
where $I_{K}$ is the isoperimetric profile of $\mathbb{H}^{n}_{K}$. Hence either $\overline{N}=0$ and the first item holds, or $\overline{N}\geq 1$. In the latter case, we want to show that $\overline{N}=1$, completing the proof of the second item. By coarea formula we can fix a sequence $\rho_{i}\nearrow+\infty$ such that $P(E\cap B_{\rho_{i}}(o))\leq P(E)+1/i$. For any $i$ we find balls $B_{s_{i}}(q_{i})\subset M\setminus B_{\rho_{i}+1}(o)$ such that $|B_{s_{i}}(q_{i})|=V-|E\cap B_{\rho_{i}}(o)|$ and such that $B_{s_{i}}(q_{i})$ converges to a ball $B\subset\mathbb{H}^{n}_{K}$ with $\lim_{i}P(B_{s_{i}}(q_{i}))=P(B)$ and $|B|=V-|E|$. If by contradiction $\overline{N}>1$, since the isoperimetric profile $I_{K}$ is a strictly subadditive function, we get
@@ 3. Local inverse mean curvature flow
We start by recalling the definition of weak inverse mean curvature flow (IMCF) as introduced in [HuiskenIlmanen].
[…]
Let $(M,g)$ be a smooth Riemannian manifold. Given a precompact $K\subset M$, a locally Lipschitz function $u:M\to\mathbb{R}$, and a set of locally finite perimeter $E$, define
@@
[…]
By virtue of [HuiskenIlmanen, Lemma 1.1], a function $u\in\mathrm{Lip}_{\mathrm{loc}}(\Omega)$ is a weak solution the IMCF in $\Omega$ if and only if
@@
The aim of this section is to show Theorem 3.1, stating that on every punctured ball $B$ centered at $o$ on a smooth complete Riemannian manifold one can define a weak IMCF that is bounded from below explicitly in terms of constants that will nicely behave on metrics $C^{0}$-close to the flat one. We also gather useful properties of this flow in Section 3.1, and Section 3.2.
[…]
@@
Indeed, one implication readily comes from the very definition in (2.1), while the other comes from an application of the coarea formula in Theorem 2.1. The latter implication is classical and dates back at least to works of Federer–Fleming and Maz’ya in the 60s, see, e.g., [FedererFleming, page 488].
We now provide solutions to the weak IMCF in punctured balls $B_{R}(o)\setminus\{o\}$. The weak IMCF is obtained in the limit, as $p\to 1^{+}$, of functions $w_{p}^{R}:=-(p-1)\log G_{p}^{R}$, where $G_{p}^{R}$, for $p\in(1,n)$ denotes the $p$-harmonic Green function on $B_{R}(o)$ with Dirichlet boundary conditions. Namely, $G_{p}^{R}$ is the solution to
{$G_{p}^{R}=\left\lvert\mathbb{S}^{n-1}\right\rvert\left(\frac{n-p}{p-1}\right)^{p-1}\delta_{o}&\text{on }B_{R}(o),\\ \phantom{-\Delta_{p}}G_{p}^{R}=0&\text{on }\partial B_{R}(o),\end{cases}$ (3.3)
where $\delta_{o}$ is the Dirac delta supported at $o$, and $\left\lvert\mathbb{S}^{n-1}\right\rvert$ is the measure of the $(n-1)$-dimensional unit sphere. With the above choice of normalization, it follows from the blow-up procedure leading to [kichenassamy-veron, Theorem 1.1] that
@@
as $r(x)\to 0$, where we let $r(x)\vcentcolon=\mathsf{d}(o,x)$. We recall that the relative $p$-capacity of a compact $K\Subset B_{R}(o)$ is defined as
@@
The following lemma is well known, and consists essentially in [holopainen-thesis, Lemma 3.8].
[…]
@@
In this proof, let $G_{p}:=G_{p}^{R}$, and $w_{p}:=w_{p}^{R}$ for the ease of notation. It is well-known that the $p$-capacity defined in (3.5) is attained computing the $L^{p}$-norm of the gradient of the $p$-harmonic function with Dirichlet data equal to $1$ on $\partial K$, and equal to 00 on $\partial B_{R}(o)$ (see [heinonen-kilpelainen-martio] for a thorough account on nonlinear potential theory). In particular, one gets, for every $t\in(0,+\infty)$,
@@
On the other hand, a straightforward application of the divergence theorem combined with the $p$-harmonicity of $G_{p}$ (see e.g. the computations in [benatti2022minkowski, Proposition 2.8]) yields that $\int_{\{G_{p}=s\}}\left\lvert\nabla G_{p}\right\rvert^{p-1}$ attains the same value for almost every $s\in(0,+\infty)$. Such constant is computed using (3.4) as
@@
$\begin{split}C_{\mathrm{cov}}(B_{\rho}(o))\vcentcolon=\min\Big{\{}N\in\mathbb{N}\ :\ &\text{$A_{3r/4,5r/4}(o)$ is covered by $N$ open balls of radius $r/2$}\\ &\text{with centers in $A_{3r/4,5r/4}(o)$ for any $0<r\leq\rho$}\Big{\}}.\end{split}$ (3.11)
@@
[…]
If $(M,g)$ in Theorem 3.1 supports a $(1,1^{*})$-Sobolev inequality, then the existence of a global proper weak IMCF issuing from a point follows from a careful modification of the proof of [KaiXu]. We mention that in the very recent [XuPhDThesis], the existence of a global solution to the weak IMCF together with the quantitative estimate (3.13) has been proven under the sole assumption that the ambient manifold supports a $(1,1^{*})$-Sobolev inequality (see [XuPhDThesis, Theorem B]).
[…]
Theorem 3.1 is analogous to the result claimed in [mari_flowlaplaceapproximationnew_2022, Theorem 1.7], and it might be argued that Theorem 3.1 follows from some adaptation of such a result. However,[mari_flowlaplaceapproximationnew_2022, Theorem 1.7] relies on the quantitative estimate claimed in [mari_flowlaplaceapproximationnew_2022, Theorem 3.6], whose proof seems to contain a gap.
The proof of Theorem 3.1 requires some preliminary steps. The following result is a direct consequence of [kotschwar_localgradientestimatesharmonic_2009, Theorem 1.1].
@@
The constant $C(K)$ in Theorem 3.2 depends on sectional curvature bounds for $g$ on $K$. The following observation will be useful also for estimating the asymptotic behavior of the Hawking mass at $o$ along the IMCF, see Section 3.1 below.
[…]
@@
Indeed, it suffices to apply [kotschwar_localgradientestimatesharmonic_2009, Equation (1.5)] on balls $B(x,\mathsf{d}(o,x)/2)$. We remark that both $\eta,\zeta$ can be chosen to be uniform with respect to $p\to 1^{+}$.
@@
It is now well-known that capacities can be estimated exploiting isoperimetric inequalities. More precisely, setting $C=1/C_{1,{\rm Sob}}(B_{2R}(o))$, we can apply [grigoryan-isocap, Eq. (7)] to get
$\begin{split}\mathrm{Cap}_{p}(\overline{B}_{r}(o),B_{2R}(o))&\geq\left(\int_{\left\lvert B_{r}(o)\right\rvert}^{\left\lvert B_{2R}(o)\right\rvert}\frac{1}{Cv^{\frac{p(n-1)}{n(p-1)}}}\right)^{1-p}\\ &=\left(C\frac{(n-p)}{n(p-1)}\right)^{p-1}\left\lvert B_{r}(o)\right\rvert^{\frac{n-p}{n}}\left[1-\left(\frac{\left\lvert B_{r}(o)\right\rvert}{\left\lvert B_{2R}(o)\right\rvert}\right)^{\frac{n-p}{n(p-1)}}\right]^{1-p}\\ &\geq\left(C\frac{(n-p)}{n(p-1)}\right)^{p-1}\left\lvert B_{r}(o)\right\rvert^{\frac{n-p}{n}}.\end{split}$
@@
Let us consider now $y\in\partial B_{r}(o)$ for $r\in(0,\tfrac{3}{2}R)$. We aim at estimating $w_{p}(y)$ combining (3.18) with a Harnack inequality. The Harnack inequality for positive $p$-harmonic functions [salvatori-rigoli-vignati, Theorem 1.28] applied to $G_{p}$ reads
@@
for any $x,y\in B_{4s/5}(z)$ such that $B_{s}(z)\subset\subset B_{2R}(o)\setminus\{o\}$, and the Harnack constant $C_{H}(B_{s}(z))$ can be estimated from above in terms of $n,C_{P}(B_{2R}(o)),C_{A}(B_{2R}(o))$ and $C_{p,\mathrm{Sob}}(B_{2R}(o))$. The interested reader might consult [mari_flowlaplaceapproximationnew_2022, Theorem 3.4 and Remark 3.5] for the full computations leading to the explicit constant in the Harnack inequality. Since $C_{p,\mathrm{Sob}}(B_{2R}(o))\to C_{1,\mathrm{Sob}}(B_{2R}(o))$ as $p\to 1^{+}$, then
@@
In order to pass to the limit $w_{p}=-(p-1)\log G_{p}$ as $p\to 1^{+}$ on compact sets $K\subset B_{2R}(o)\setminus\{o\}$, we need also some uniform upper bound on $w_{p}$. This is a well-known consequence of the Laplacian comparison theorem and of the asymptotics of the Green function at the pole in (3.4), see [Kura, Theorem 1.2].
[…]
@@
Proof of Theorem 3.1..
Within this proof let $G_{p}^{2R}=:G_{p}$, and $w_{p}^{2R}=:w_{p}$ for the ease of notation. Let $K\subset B_{2R}(o)\setminus\{o\}$ be compact. We first show that $w_{p}$ is equibounded on $K$ as $p\to 1^{+}$. Indeed by (3.16) we know that $w_{p}\geq(n-p)\log(r(x))-C$ on $B_{3R/2}(o)\setminus\{o\}$ for $C$ as in Theorem 3.3. On the other hand, taking $\overline{r}\in(0,2R)$ such that $K\subset B_{\overline{r}}(o)\setminus\{o\}\subset\subset B_{2R}(o)$, if $\mathrm{Ric}\geq-(n-1)a$ on $B_{2R}(o)$ for some $a>0$, by (3.22) we have that for all $x\in B_{\overline{r}}(o)\setminus\{o\}$ there holds
@@
We claim that $w$ satisfies the weak formulation of the IMCF on $B_{2R}(o)\setminus\{o\}$. Indeed, arguing as in [MoserIMCF, Equation (9)], one gets that $|\nabla w_{p}|^{p}\mathcal{H}^{n}$ converges to $|\nabla w|\mathcal{H}^{n}$ in duality with bounded continuous functions on compact subsets contained in $B_{2R}(o)\setminus\{o\}$, along the suitable sequence $p_{i}\to 1^{+}$ for which $w_{p_{i}}$ converges. Using again [MoserIMCF, Equation (9)] this is enough to conclude that $w$ is a weak solution of the IMCF on $B_{2R}(o)\setminus\{o\}$.
@@
Therefore we proved items (1), (2) and (3) of the statement. It remains to show that $\{w\leq(n-1)\log r-C-1\}\subset B_{r}(o)$ for every $r\leq R$. Fix $r\in(0,R]$. By Theorem 3.3 there exists

### v3

@@ 2.3. Technical Lemmas on $C^{0}$-Riemannian manifolds
Combining again the fact that $M$ is $C^{0}_{\rm loc}$-asymptotic to $\mathbb{H}^{n}_{K}$ with item (1) in Section˜2.3 we get that for any $R>0$ there exists a constant $C_{R}>0$ such that $C_{R}^{-1}r^{n}\leq|B_{r}(x)|\leq C_{R}r^{n}$ for any $x\in M$ and $r\in(0,R)$. Hence recalling that $(M,\mathsf{d},\mathcal{H}^{n})$ is also PI by Section˜2.3, then it is well-known that there holds a relative isoperimetric inequality in balls of $M$, see [34, Theorem 5.1] and [3, Remark 4.4]. Since also $\inf_{x\in M}|B_{1}(x)|>0$ thanks to the asymptotic assumption, it is readily checked that the proof of [5, Lemma 2.10] can be repeated in our setting, yielding that: for any $R>0$ there exists $\widetilde{C}=\widetilde{C}(R)>0$ such that for any $E\subset M$ with $|E|\in(0,+\infty)$ there exists $x_{E}$ such that
@@
Fix now $\overline{V}\in(0,+\infty)$ and $\eta\in(0,\overline{V})$. Let also $V_{0}\in(\max\{\overline{V}-1,\eta\},\overline{V}+1)$ and for any $\varepsilon>0$ let $E\subset M$ be a bounded set such that $|E|=V_{0}$ and $P(E)\leq I(V_{0})+\varepsilon$ (it is readily checked that, arguing as in [4, Lemma 2.17], the isoperimetric profile is achieved by bounded sets).
@@
The next theorem is based on a concentration-compactness argument that has been used several times in the literature, applied to the study of the isoperimetric problem in noncompact manifolds. In the smooth setting it has been first obtained in [59, Theorem 2]. See also [65, Theorem 4.48] and references therein for a complete account. Results analogous to Theorem˜2.16 have been worked out also in the setting of nonsmooth spaces with bounds below on the curvature, see [4, Theorem 4.6] and [5, Theorem 3.3 & Theorem 1.1].
@@
Since the proof of Theorem˜2.16 is standard and follows closely the strategy of [4, 5] we just sketch it.
At first, one gets the analogue of Ritoré–Rosales’ result [5, Theorem 3.3] in the setting of Theorem˜2.16. Indeed, the proof of [5, Theorem 3.3] uses the coarea formula, the precompactness of $BV$ in $L^{1}_{\mathrm{loc}}$, and the existence, around every point $p\in M$, of a one-parameter family $\{\mathcal{B}_{p,r}\}_{r\in(0,\varepsilon)}$ of sets such that $r\mapsto|\mathcal{B}_{p,r}|$ is continuous and vanishing as $r\to 0$, and $P(\mathcal{B}_{p,r})\to 0$ as $r\to 0$. The first two come from Section˜2.3, and Theorem˜2.7, while for the last one it suffices to take pre-images of small Euclidean balls under the map in the second item of Section˜2.1, as it has been done in the proof of Section˜2.3. Once this is done, one follows verbatim the proof of [4, Theorem 4.6], which additionally needs that $(M,\mathsf{d},\mathcal{H}^{n})$ is PI, and $|B_{r}(p)|\geq\zeta r^{3}$ for every $r\in(0,R]$, and every $p\in M$, where $\zeta,R>0$ are constants depending on $M$: these two properties come from Section˜2.3 and from its proof (compare also with item (1) in Section˜2.3).
By Section˜2.3, the metric measure space $(M,\mathsf{d},\mathcal{H}^{n})$ is PI, hence it satisfies a relative local isoperimetric inequality (as already explained in the proof of Section˜2.3). We can thus directly follow [4, Theorem 4.6] until (4.20) there, and then jump to Step 5 there. One gets the following. There is $\overline{N}\in\mathbb{N}\cup\{+\infty\}$ and radii $R_{i}\to+\infty$, $T_{i,j}\to_{i}+\infty$ for $1\leq j<\overline{N}+1$, and there are mutually (with respect to $j$) diverging $p_{i,j}\in M\setminus B_{R_{i}}(o)$ such that
$E_{i}^{c}\xrightarrow[i]{}E\text{ in $L^{1}(M)$,}\qquad(E_{i}\setminus B_{R_{i}}(o))\cap B_{T_{i,j}}(p_{i,j})\xrightarrow[i]{}B_{j}\text{ in $L^{1}(M)$,}$
where $B_{j}$ is a ball in $\mathbb{H}^{n}_{K}$ for any $j<\overline{N}+1$, and $E$ is an isoperimetric set in $M$. The convergence in $L^{1}$ to the $B_{j}$’s has to be intended as in the statement of Section˜2.3, after the composition with biLipschitz embeddings. Moreover one has
@@
where $I_{K}$ is the isoperimetric profile of $\mathbb{H}^{n}_{K}$. Hence either $\overline{N}=0$ and the first item holds, or $\overline{N}\geq 1$. In the latter case, we want to show that $\overline{N}=1$, completing the proof of the second item. By coarea formula we can fix a sequence $\rho_{i}\nearrow+\infty$ such that $P(E\cap B_{\rho_{i}}(o))\leq P(E)+1/i$. For any $i$ we find balls $B_{s_{i}}(q_{i})\subset M\setminus B_{\rho_{i}+1}(o)$ such that $|B_{s_{i}}(q_{i})|=V-|E\cap B_{\rho_{i}}(o)|$ and such that $B_{s_{i}}(q_{i})$ converges to a ball $B\subset\mathbb{H}^{n}_{K}$ with $\lim_{i}P(B_{s_{i}}(q_{i}))=P(B)$ and |B|=V|E||B|=V-|E|. If by contradiction $\overline{N}>1$, since the isoperimetric profile $I_{K}$ is a strictly subadditive function, we get
@@ 3. Local inverse mean curvature flow
We start by recalling the definition of weak inverse mean curvature flow (IMCF) as introduced in [38].
Definition 3.1 (Weak IMCF - Level set formulation).
Let $(M,g)$ be a smooth complete Riemannian manifold. Given a precompact $K\subset M$, a locally Lipschitz function $u:M\to\mathbb{R}$, and a set of locally finite perimeter $E$, define
@@
Remark 3.2.
By virtue of [38, Lemma 1.1], a function $u\in\mathrm{Lip}_{\mathrm{loc}}(\Omega)$ is a weak solution the IMCF in $\Omega$ if and only if
@@
The aim of this section is to show Theorem˜3.6, stating that on every punctured ball $B$ centered at $o$ on a smooth complete Riemannian manifold one can define a weak IMCF that is bounded from below explicitly in terms of constants that will nicely behave on metrics $C^{0}$-close to the flat one. We also gather useful properties of this flow in Section˜3.1, and Section˜3.2.
Definition 3.3.
@@
Indeed, one implication readily comes from the very definition in (2.1), while the other comes from an application of the coarea formula in Theorem˜2.7. The latter implication is classical and dates back at least to works of Federer–Fleming and Maz’ya in the 60s, see, e.g., [29, page 488].
Let $(M,g)$ be a smooth complete Riemannian manifold. We now provide solutions to the weak IMCF in punctured balls $B_{R}(o)\setminus\{o\}$. The weak IMCF is obtained in the limit, as $p\to 1^{+}$, of functions $w_{p}^{R}:=-(p-1)\log G_{p}^{R}$, where $G_{p}^{R}$, for $p\in(1,n)$ denotes the $p$-harmonic Green function on $B_{R}(o)$ with Dirichlet boundary conditions. Namely, $G_{p}^{R}$ is the solution to
[…]
where $\delta_{o}$ is the Dirac delta supported at $o$, and $\left\lvert\mathbb{S}^{n-1}\right\rvert$ is the measure of the (n1)(n-1)-dimensional unit sphere. With the above choice of normalization, it follows from the blow-up procedure leading to [44, Theorem 1.1] that
@@
as $r(x)\to 0$, where we let $r(x)\vcentcolon=\mathsf{d}(o,x)$. A full proof of (3.4) can be found in the recent [13, Theorem 2.1]. We recall that the relative $p$-capacity of a compact $K\Subset B_{R}(o)$ is defined as
@@
The following lemma is well known, and consists essentially in [37, Lemma 3.8].
Lemma 3.5.
@@
In this proof, let $G_{p}:=G_{p}^{R}$, and $w_{p}:=w_{p}^{R}$ for the ease of notation. It is well-known that the $p$-capacity defined in (3.5) is attained computing the $L^{p}$-norm of the gradient of the $p$-harmonic function with Dirichlet data equal to $1$ on $\partial K$, and equal to 0 on $\partial B_{R}(o)$ (see [36] for a thorough account on nonlinear potential theory). In particular, one gets, for every $t\in(0,+\infty)$,
@@
On the other hand, a straightforward application of the divergence theorem combined with the $p$-harmonicity of $G_{p}$ (see e.g. the computations in [10, Proposition 2.8]) yields that $\int_{\{G_{p}=s\}}\left\lvert\nabla G_{p}\right\rvert^{p-1}$ attains the same value for almost every $s\in(0,+\infty)$. Such constant is computed using (3.4) as
@@
[…]
@@
Remark 3.7.
If $(M,g)$ in Theorem˜3.6 supports a $(1,1^{*})$-Sobolev inequality, then the existence of a global proper weak IMCF issuing from a point follows from a careful modification of the proof of [73]. We mention that in the very recent [74], the existence of a global solution to the weak IMCF together with the quantitative estimate (3.13) has been proven under the sole assumption that the ambient manifold supports a $(1,1^{*})$-Sobolev inequality (see [74, Theorem B]).
Remark 3.8.
Theorem˜3.6 is analogous to the result claimed in [53, Theorem 1.7], and it might be argued that Theorem˜3.6 follows from some adaptation of such a result. However,[53, Theorem 1.7] relies on the quantitative estimate claimed in [53, Theorem 3.6], whose proof seems to contain a gap.
The proof of Theorem˜3.6 requires some preliminary steps. The following result is a direct consequence of [47, Theorem 1.1].
@@
The constant $C(K)$ in Theorem˜3.9 depends on sectional curvature bounds for $g$ on $K$. The following observation will be useful also for estimating the asymptotic behavior of the Hawking mass at $o$ along the IMCF, see Section˜3.1 below.
Remark 3.10.
@@
Indeed, it suffices to apply [47, Equation (1.5)] on balls $B(x,\mathsf{d}(o,x)/2)$. We remark that both $\eta,\zeta$ can be chosen to be uniform with respect to $p\to 1^{+}$.
@@
It is now well-known that capacities can be estimated exploiting isoperimetric inequalities. More precisely, setting $C=1/C_{1,{\rm Sob}}(B_{2R}(o))$, we can apply [32, Eq. (1.7) p. 141] to get
$\begin{split}\mathrm{Cap}_{p}(\overline{B}_{r}(o),B_{2R}(o))&\geq\left(\int_{\left\lvert B_{r}(o)\right\rvert}^{\left\lvert B_{2R}(o)\right\rvert}\frac{1}{Cv^{\frac{p(n-1)}{n(p-1)}}}\,\mathrm{d}v\right)^{1-p}\\ &=\left(C\frac{(n-p)}{n(p-1)}\right)^{p-1}\left\lvert B_{r}(o)\right\rvert^{\frac{n-p}{n}}\left[1-\left(\frac{\left\lvert B_{r}(o)\right\rvert}{\left\lvert B_{2R}(o)\right\rvert}\right)^{\frac{n-p}{n(p-1)}}\right]^{1-p}\\ &\geq\left(C\frac{(n-p)}{n(p-1)}\right)^{p-1}\left\lvert B_{r}(o)\right\rvert^{\frac{n-p}{n}}.\end{split}$
@@
Let us consider now $y\in\partial B_{r}(o)$ for $r\in(0,\tfrac{3}{2}R)$. We aim at estimating $w_{p}(y)$ combining (3.18) with a Harnack inequality. The Harnack inequality for positive $p$-harmonic functions [64, Theorem 1.28] applied to $G_{p}$ reads
@@
for any $x,y\in B_{4s/5}(z)$ such that $B_{s}(z)\subset\subset B_{2R}(o)\setminus\{o\}$, and the Harnack constant $C_{H}(B_{s}(z))$ can be estimated from above in terms of $n,C_{P}(B_{2R}(o)),C_{A}(B_{2R}(o))$ and $C_{p,\mathrm{Sob}}(B_{2R}(o))$. The interested reader might consult [53, Theorem 3.4 and Remark 3.5] for the full computations leading to the explicit constant in the Harnack inequality. Since $C_{p,\mathrm{Sob}}(B_{2R}(o))\to C_{1,\mathrm{Sob}}(B_{2R}(o))$ as $p\to 1^{+}$, then
@@
In order to pass to the limit $w_{p}=-(p-1)\log G_{p}$ as $p\to 1^{+}$ on compact sets $K\subset B_{2R}(o)\setminus\{o\}$, we need also some uniform upper bound on $w_{p}$. This is a well-known consequence of the Laplacian comparison theorem and of the asymptotics of the Green function at the pole in (3.4). A proof of the following result can be found in [48, Theorem 2].
Proposition 3.12.
@@
Proof of Theorem˜3.6..
Within this proof let $G_{p}^{2R}=:G_{p}$, and $w_{p}^{2R}=:w_{p}$ for the ease of notation. Let $K\subset B_{2R}(o)\setminus\{o\}$ be compact. We first show that $w_{p}$ is equibounded on $K$ as $p\to 1^{+}$. Indeed by (3.16) we know that $w_{p}\geq(n-p)\log(r(x))-C$ on $B_{3R/2}(o)\setminus\{o\}$ for $C$ as in Theorem˜3.11. On the other hand, taking $\overline{r}\in(0,2R)$ such that $K\subset B_{\overline{r}}(o)\setminus\{o\}\subset\subset B_{2R}(o)$, if $\mathrm{Ric}\geq-(n-1)a$ on $B_{2R}(o)$ for some $a>0$, by (3.22) we have that for all $x\in B_{\overline{r}}(o)\setminus\{o\}$ there holds
@@
We claim that $w$ satisfies the weak formulation of the IMCF on $B_{2R}(o)\setminus\{o\}$. Indeed, arguing as in [56, Equation (9)], one gets that $|\nabla w_{p}|^{p}\mathcal{H}^{n}$ converges to $|\nabla w|\mathcal{H}^{n}$ in duality with bounded continuous functions on compact subsets contained in $B_{2R}(o)\setminus\{o\}$, along the suitable sequence $p_{i}\to 1^{+}$ for which $w_{p_{i}}$ converges. Using again [56, Equation (9)] this is enough to conclude that $w$ is a weak solution of the IMCF on $B_{2R}(o)\setminus\{o\}$.
@@
Therefore we proved items (1), (2) and (3) of the statement. It remains to show that $\{w\leq(n-1)\log r-C-1\}\subset B_{r}(o)$ for every $r\leq R$. Fix $r\in(0,R]$. By Theorem˜3.11 there exists